OpenStudy (anonymous):

Medal and fanning who answers correctly first. What is the solution to the equation? 4x - 26 - 3x= 16 - 20

OpenStudy (babylove18):

4x-26-3x=16-20 add like terms x-26 = -4 add 26 to both sides x=22

OpenStudy (anonymous):

he is right

OpenStudy (anonymous):

ty @babylove18

OpenStudy (babylove18):

im a female ! and your welcome @ItzHannahx

OpenStudy (anonymous):

sorry about that